OpenStudy (anonymous):

solve 5/8x+1/16x=5/16+x

OpenStudy (anonymous):

10/16x+ 1/16x=5/16+x 11/16x=5/16+x -5/16x=5/16 -5x=5 x=-1

jhonyy9 (jhonyy9):

16+x is denominator ??? - or just 16 is ?

OpenStudy (anonymous):

16x